What is the change in entropy in converting 1.0 kg of ice at 0.0-degree celsius to 1.0 kg water at 0.0 degrees celsius?


1
Expert's answer
2021-02-17T11:00:25-0500

"\lambda=340~\frac{kJ}{kg},"

"\Delta S=\frac QT,"

"\Delta S=\frac{340\cdot 10^3\cdot 1}{273}=1245~\frac JK."